OpenStudy (anonymous):

((2+i)/(3-i))+((2-i)/(4i)) explanation please

OpenStudy (anonymous):

do u know that these are complex numbers eg 2 + i

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u need to rationalise the denominators by multiplying them by their complements the complement of 3-i is 3+i (changed sign) so we have (2+i)((3+i)/(3-i)(3+i) + (2-i)4i / 4i(4i) (6+2i+3i -1)/ 9+1 + 8i +4/-16 (5 +5i )/ 10 - (8i+4)16

OpenStudy (anonymous):

what do i do after that?

OpenStudy (anonymous):

i suppose we should write it as one fraction (1+i) / 2 - (2i+1) / 4 = (2+2i )/ 4 - (2i+1) / 4 = 1/4

OpenStudy (anonymous):

do i need a common denominator?

OpenStudy (anonymous):

wow - that neat- if i'm right of couse

OpenStudy (anonymous):

you are i checked it on my calculator but i didnt know how to get it.... and i factor right?

OpenStudy (anonymous):

sorry i dont know exactly what you mean (i'm a brit see!!)

OpenStudy (anonymous):

nevermind but thanks

OpenStudy (anonymous):

i understand it now

OpenStudy (anonymous):

good